OpenStudy (anonymous):

2. A block of material has a volume of 2 m3. It has a mass of 1 × 104 kg. Calculate the density of the material sample and express your results in g/cm3 with appropriate scientific notation.

OpenStudy (anonymous):

Density = mass/volume plug in the numbers

OpenStudy (joannablackwelder):

First convert 1x10^4 kg into g (1000 g = 1 kg) and 2 m^3 into cm^3 (1 m = 100 cm => 1 m^3 = 100^3 cm^3). then plug in the numbers. into the formula.